OpenStudy (anonymous):

Solve the equations: x^2+y^2=169, x-y-13=0. . . . . ANSWER = {(13,0),(0,-13)}

OpenStudy (anonymous):

x = y + 13

OpenStudy (anonymous):

\[(y + 13)^2 = y^2 + 26y +169\]

hartnn (hartnn):

write x^2+y^2 = (x-y)^2 +2xy

OpenStudy (anonymous):

so: \[2y^2 + 26 +169 = 169\]

OpenStudy (anonymous):

\[2y(y+13) = 0\]

OpenStudy (anonymous):

forgot a y two posts up

OpenStudy (anonymous):

y = 0, or 13

OpenStudy (anonymous):

the rest is trivial

OpenStudy (anonymous):

negative 13

OpenStudy (anonymous):

typos :(

OpenStudy (anonymous):

how did 2y(y+13)=0

OpenStudy (anonymous):

you cancel 169, and factor out a 2y

OpenStudy (anonymous):

i am not gettin it

OpenStudy (anonymous):

2y^2+26+169=169 HOW!!!!!

OpenStudy (anonymous):

\[(y+13)^2 + y^2 = 169\] \[(y^2 + 26y + 169) + y^2 = 169\] \[2y^2 + 26 + 169 = 169\]

OpenStudy (anonymous):

what about the values of x? how will i get those

OpenStudy (anonymous):

then you just use the second equation

OpenStudy (anonymous):

if y = 0: \[x-(0)-13 = 0\]

OpenStudy (anonymous):

then x = 13

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so the ordered pair is (13,0)

OpenStudy (anonymous):

ok done@ pretty easy method of solving thanks

OpenStudy (anonymous):

Yeah that's right @MrMoose It's correct

OpenStudy (anonymous):

You are welcome